Consider the following.

a. What is the duration of a four-year Treasury bond with a 7 percent semiannual coupon selling at par?
b. What is the duration of a three-year Treasury bond with a 7 percent semiannual coupon selling at par?
c. What is the duration of a two-year Treasury bond with a 7 percent semiannual coupon selling at par?
Answer #1

a

Period Cash Flow Discounting factor PV Cash Flow Duration Calc
0 ($1,000.00) =(1+YTM/number of coupon payments in the year)^period =cashflow/discounting factor =PV cashflow*period
1             35.00                                                             1.04                    33.82                  33.82
2             35.00                                                             1.07                    32.67                  65.35
3             35.00                                                             1.11                    31.57                  94.70
4             35.00                                                             1.15                    30.50                122.00
5             35.00                                                             1.19                    29.47                147.35
6             35.00                                                             1.23                    28.47                170.84
7             35.00                                                             1.27                    27.51                192.57
8       1,035.00                                                             1.32                  785.99              6,287.93
      Total              7,114.54
Macaulay duration =(∑ Duration calc)/(bond price*number of coupon per year)
=7114.54/(1000*2)
=3.557272

b

Cash Flow Discounting factor PV Cash Flow Duration Calc
($1,000.00) =(1+YTM/number of coupon payments in the year)^period =cashflow/discounting factor =PV cashflow*period
            35.00                                                             1.04                    33.82                  33.82
            35.00                                                             1.07                    32.67                  65.35
            35.00                                                             1.11                    31.57                  94.70
            35.00                                                             1.15                    30.50                122.00
            35.00                                                             1.19                    29.47                147.35
      1,035.00                                                             1.23                  841.97              5,051.84
      Total              5,515.05
Macaulay duration =(∑ Duration calc)/(bond price*number of coupon per year)
=5515.05/(1000*2)
=2.757526

c

Period Cash Flow Discounting factor PV Cash Flow Duration Calc
0 ($1,000.00) =(1+YTM/number of coupon payments in the year)^period =cashflow/discounting factor =PV cashflow*period
1             35.00                                                             1.04                    33.82                  33.82
2             35.00                                                             1.07                    32.67                  65.35
3             35.00                                                             1.11                    31.57                  94.70
4       1,035.00                                                             1.15                  901.94              3,607.77
      Total              3,801.64
Macaulay duration =(∑ Duration calc)/(bond price*number of coupon per year)
=3801.64/(1000*2)
=1.900818